3p=4-9p
+9p=4=9p
12p=4
12p/4=3p

Found 2 solutions by rfer, yonelleb:
Answer by rfer(16322)

3p=4-9p
12p=4
p=4/12
p=1/3

Answer by yonelleb(1)